# SAMSA: Efficient Transformer for Many Data Modalities

### v1

@@
Abstract.
The versatility of self-attention mechanism earned transformers great success in almost all data modalities, with limitations on the quadratic complexity and difficulty of training. Efficient transformers, on the other hand, often rely on clever data-modality-dependent construction to get over the quadratic complexity of transformers. This greatly hinders their applications on different data modalities, which is one of the pillars of contemporary foundational modeling. In this paper, we lay the groundwork for efficient foundational modeling by proposing SAMSA - SAMpling-Self-Attention, a context-aware linear complexity self-attention mechanism that works well on multiple data modalities. Our mechanism is based on a differentiable sampling without replacement method we discovered. This enables the self-attention module to attend to the most important token set, where the importance is defined by data. Moreover, as differentiability is not needed in inference, the sparse formulation of our method costs little time overhead, further lowering computational costs. In short, SAMSA achieved competitive or even SOTA results on many benchmarks, while being faster in inference, compared to other very specialized models. Against full self-attention, real inference time significantly decreases while performance ranges from negligible degradation to outperformance. We release our source code in the repository: https://github.com/HySonLab/SAMSA.
differentiable algorithm,efficient transformer,deep learning
Correspondence to Truong Son Hy at thy@uab.edu
ccs: Computing methodologies Neural networksccs: Computing methodologies Supervised learning
1. Introduction
@@ 2. Related works
The scalability of transformers is hindered by the quadratic complexity of self-attention modules. Hence, numerous works have delved into sparsity (Beltagy et al., 2020; Ainslie et al., 2020; Zaheer et al., 2020; Shirzad et al., 2023), low-rank projection (Wang et al., 2020), hashing-based method (Kitaev et al., 2020), and kernel-based methods (Choromanski et al., 2021; Katharopoulos et al., 2020) to sacrifice the expressivity of transformers for a sub-quadratic construction. Our review of contemporary efficient transformers is influenced by this survey by (Tay et al., 2022). Linformer (Wang et al., 2020) is based on the low-rank assumption of self-attention matrices, realized via projection matrices. As one of the earliest works in efficient transformers, it has flaws in both scalability and fixed sequence length. Subsequent work like Linear Transformer (Katharopoulos et al., 2020) redesigns the similarity kernel previously as a softmax of dot-product score (or any other kernels) to one made of a linear probability cage. The modification frees them from the recomputation of the multiplication of key-value matrices, making attention computational cost $O(N)$, where $N$ refers to the number of tokens. Without softmax non-linearity, the rank of those attention matrices is bounded by the number of dimensions of tokens, which loses much expressivity of transformer models. Methods like (Beltagy et al., 2020) utilize three choices of fixed-sparsity-patterns: sliding window, dilated sliding window, and global+sliding window. Similarly, (Zaheer et al., 2020) uses a combination of different sparse patterns at one time: random attention, window attention, and global attention. On graphs, (Shirzad et al., 2023) also combines different sparsity patterns but with a constraint on the random-generated one: it has to be an expander graph. The hard-coded sparsity effectively linearized the complexity of transformers and is empirically fast. However, intuitively, no pattern or finite set of patterns fits all problems; therefore, there is also a line of work dedicated to a differentiable sparsity. (Roy et al., 2021) drew a relationship from MIPS problem (Maximum Inner Product Search) and the Nearest-Neighbor Search algorithm, when both the query and key vectors are unit vectors. With their online k-means algorithm, their method attends each query vector to every key vector belonging to the same cluster, thus, bringing the complexity down to $O(n^{1.5})$, with $n$ is the number of tokens. Practically, however, their method is not fast as it needs a k-means construction and extensive specialized implementation for sparse operators. Another perspective on efficient transformers is making better implementation of vanilla transformers (Rabe and Staats, 2021; Dao et al., 2022; Dao, 2024). (Rabe and Staats, 2021) proposed a neat trick for attention mechanism: chunking. While the mechanism is simple, they have reduced the memory cost from quadratic to sublinear, effectively lifting the computation constraint of GPU memory. (Dao et al., 2022) both reduces the frequency of data transfer between HBM and SRAM within the GPU and derives a memory efficient backward pass. This allows quadratic attention scales to the sequence length of 64000 and beyond. (Dao, 2024) is an incremental work from (Dao et al., 2022) with better work partition and reduction in non-matmul FLOPS.
@@
(1) $H(x)=\begin{cases}0&\text{if }x<0\\ 1&\text{if }x\geq 0\end{cases}$
@@
has gradient undefined at $x=0$ and zero everywhere else. Since continuous relaxation of discrete operators and algorithms is a complex study and the sampling operator is the core of our efficient transformer formulation, we focus our literature review on differentiable sampling. To continuously relax the sampling operator, the previous literature was developed based on the relaxation of choose, sort, and top-k algorithms (Jang et al., 2017; Maddison et al., 2017; Kool et al., 2019; Xie et al., 2020; Blondel et al., 2020; Petersen et al., 2021). (Jang et al., 2017; Maddison et al., 2017) is the foundational work on parameterization of discrete categorical variable. They introduces the Gumbel-Softmax distribution that approximate categorical distribution exceptionally well. There are some differences between the two works: (Jang et al., 2017) introduces straight through estimator via application of path derivative (Bengio et al., 2013) and (Maddison et al., 2017) takes account of density in their relaxed objective function. From differentiable categorical variable, differentiable sampling with replacement can be obtained trivially by computing these stochastic categorical variable $k$-times, where $k$ is the number of samples. Differentiable sampling without replacement, however, is much trickier and is an active area of research. (Kool et al., 2019) is one of the earliest work in differentiable sampling without replacement. Their process of sampling without replacement, Gumbel Top-k, samples tokens sequentially with logit normalization, applied in generative text models as stochastic beam search. (Xie et al., 2020) sees the differentiable without replacement in another perspective: optimal transport, where loss function represents how far away is the current solution from the optimal one. (Blondel et al., 2020) is an efficient estimation of sort function differentiation through optimal transport via permutahedron projections. The projection decrease quadratic computational cost to $O(n\log{}n)$ forward pass and $O(n)$ backward pass. (Petersen et al., 2021) takes a pre-existing discrete parallelizable sorting/top-k algorithm (e.g. odd-even sorting network, bitonic sorting network, selection network, …) and relaxes all of the control flows for gradient estimation.
3. Background
3.1. Transformer Architecture
@@
The self-attention mechanism is a core component of Transformer models to capture long-range dependencies. Let $\mathbf{X}=\left\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\right\}$ $\in\mathbb{R}^{n\times d}$ represent a sequence of $n$ $d$-dimensional vectors. The matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ all in $\mathbb{R}^{n\times d_{h}}$ representing query, key, and value vectors are created using linearity:
$\mathbb{Q}$ $\;=\mathbb{X}W_{Q}+\mathbb{1}_{n}b_{q}^{\top},$
$\mathbb{K}$ $\;=\mathbb{X}W_{K}+\mathbb{1}_{n}b_{k}^{\top},$
$\mathbb{V}$ $\;=\mathbb{X}W_{V}+\mathbb{1}_{n}b_{v}^{\top}.$
Then, using the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ vectors, it creates another sequence by aggregating tokens satisfying certain criteria, expressed as follows:
@@
(2) $\operatorname{Attn}(\mathbf{X})=\operatorname{SM}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V},$
where $\operatorname{SM}$ is the softmax nonlinearity. The output of the self-attention layer is propagated through a token-wise feed forward
(3) $\operatorname{FFN}(\mathbb{X})=\sigma(\mathbb{X}W_{1}+\mathbb{1}_{n}b_{1}^{\top})W_{2}+\mathbb{1}_{n}b_{2}^{\top},$
where $\sigma$ is an elementwise nonlinearity. The intrinsic problem of Transformer is the calculation of $\operatorname{SM}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{d_{h}}\right)$ in Eq.2, which calculate pair-wise attention score, leading to $O(n^{2})$ computation cost. Despite the computational complexity of transformers being quadratic, the memory complexity of transformers can be reduced to $O(n)$, shown in (Rabe and Staats, 2021; Dao et al., 2022; Dao, 2024). The computational cost of self-attention is further decreased through fewer memory operations between HBM and SRAM and better parallelization in (Dao et al., 2022; Dao, 2024). These improvements are compatible with our model and are integrated into our model via existing fused attention implementation.
3.2. Continuous Relaxation of Top-$k$ Sampling
A methodology to find the optimal sparse pattern for attention is too make such process learnable, i.e. differentiable and convergent. But evidently, top-$k$ sampling methods are not differentiable. A way to work around this is by relaxing the sampling function, represented by the sampling matrix, by an soft alternative. The criteria for choosing such a relaxation is such that there is some hyperparameter that can be tuned to approximate the hard function from the soft function with abitrary accuracy, for instance the softmax and the hardmax. The goal of this technique is to differentiablize functions without derivatives.
In the context of a top-$k$ sampling function $\mathcal{s}:\mathbb{R}^{n}\to\mathbb{R}^{k}$, these functions can be represented as a transformation $S(x)=\mathbb{P}x$ where $\mathbb{P}\in\{0,1\}^{k\times n}$ and each row of $\mathbb{P}$ contains one value 1. If $\mathbb{P}$ is fixed and independent of $x$, then $S(x)$ is differentiable and such methods are fixed sampling, or fixed sparsity patterns such as sliding window, random attention, this is not possible if we let $\mathbb{P}=\mathbb{P}(x)$. The common solution to this is approximating $\mathbb{P}$ by a softmax output, which is continuous. One can also use the soft approximation solely for the backward calculation, this is called the reparametrization trick. To examplify reparametrization trick, we present an optimization problem as a toy example.
Example 3.1.
Given an input vector $\mathbb{x}\in\mathbb{R}^{10}$ and a label $y\in\mathbb{R}^{5}$, minimize the weight sum of five values of $\mathbb{x}$ against $y$. In other words, solve the optimization problem
$\min_{\begin{subarray}{c}{i_{j},w,b}\\ {j\in[5]}\end{subarray}}\sum_{j=1}^{5}\left(x_{i_{j}}w+b-y_{j}\right)^{2}.$
To solve Example 3.1, our goal is to design a neural network which features sampling and obtains the three largest values. Let us call assume our network be the composition $f\circ g$ where $f$ be the simple neural network defined by $w$ and $b$, while $g$ is the sampling function selecting five most relevent inputs in $\mathbb{x}$. Assume that $f\circ g$ lies in the hypothesis space, which means that it can perfectly learn, i.e. find the hard threshold for three largest values of $\mathbb{x}$. In our design, the sampling function $g:\mathbb{R}^{10}\to\mathbb{R}^{5}$ should be
$g(\mathbb{x})=\mathbb{P(x)x}$
where $\mathbb{P}:\mathbb{R}^{10}\to\mathbb{R}^{5}$ is the sampling matrix that is value-aware. However, $g$ is clearly not differentiable, and, thus, cannot be trained in practice. To handle this, we create a backward soft proxy for $g$ called $g_{soft}$ where during training,
$\nabla g\triangleq\nabla g_{soft}$
but the forward process remains the same. This is the main concept of the reparametrization trick.
4. Method
4.1. Overview
@@
Figure 1. Overview of our proposed model sampling-self-attention module SAMSA. The key-value vectors are selected via top-$k$ importance score computed using tokens’ latent. The Gumbel-Sigmoid reparameterization provides gradients to guide the optimization process towards most important key-value pairs of vectors (left). The sampled key-value vectors are then fed into Flash Attention to attend to the query vectors (right).
@@
1:  Transform token matrix to query, key, and value vectors
@@
3:  $\mathbf{Q}\leftarrow\widehat{\mathbb{X}}W_{Q}+\mathbb{1}_{n}b_{q}^{\top}$
4:  $\mathbf{K}\leftarrow\widehat{\mathbb{X}}W_{K}+\mathbb{1}_{n}b_{k}^{\top}$
5:  $\mathbf{V}\leftarrow\widehat{\mathbb{X}}W_{V}+\mathbb{1}_{n}b_{v}^{\top}$
6:  Compute importance score
7:  $\mathbf{S}\leftarrow\operatorname{ImportanceScore}(\mathbf{X})+\operatorname{Gumbel}(0,1)$
8:  $\mathbb{P}\leftarrow\operatorname{SamplingMatrix}(\mathbb{S},k)$Refer to Algorithm 2}
9:  $\tilde{\mathbf{K}},\tilde{\mathbf{V}}\leftarrow\mathbb{PK},\mathbb{PV}$
10:  Compute attention score
11:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FlashAttention}(\mathbf{Q},\tilde{\mathbf{K}},\tilde{\mathbf{V}})$
12:  Feed-Forward Network
13:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FFN}(\widehat{\mathbb{X}})$
14:  Return $\mathbf{X}$
4.2. Differentiable Top-$k$ Sampling
Sampling operator definition.
We formally define the operator that samples $k$ of $n$ real $d$-dimensional vectors as a linear function $\operatorname{S}(\mathbf{X},k):\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{k\times d}=\mathbf{P}\mathbf{X}$. Here, $\mathbf{P}$ the sampling matrix is a row-wise concatenation of one-hot vectors defined as follows:
(4) $\mathbb{P}=\left[\begin{array}[]{ccccc}\operatorname{onehot}(x_{1})\\ \operatorname{onehot}(x_{2})\\ \vdots\\ \operatorname{onehot}(x_{k})\\ \end{array}\right].$
where $x_{1\dots k}$ indicates the one-hot vector represents the indices of the most important tokens.
Score-based Context-Aware Sampling.
Here, we define importance scores as the tokens’ likeliness of being picked. The importance scores of tokens are computed via a neural network: $\operatorname{NN}(\mathbf{X}):\mathbb{R}^{nm}\rightarrow\mathbb{R}^{n}$. While this method is simple, it conditions the importance score of each token on the entire set of tokens, allowing patterns with complex dependencies to emerge. For randomness, we use $\operatorname{Gumbel}$ distribution to include randomness to the computed importance scores, similar to (Jang et al., 2017). The Gumbel-Distribution PDF and importance score computation is given as follows:
$\operatorname{Gumbel}(0,1)(x)$ $=e^{-(x+e^{-x})},$
$\operatorname{ImportanceScore}(\mathbb{X})$ $=\operatorname{NN}(\mathbb{X}).$

### v2

@@
Abstract
The versatility of self-attention mechanism earned transformers great success in almost all data modalities, with limitations on the quadratic complexity and difficulty of training. Efficient transformers, on the other hand, often rely on clever data-modality-dependent construction to get over the quadratic complexity of transformers. This greatly hinders their applications on different data modalities, which is one of the pillars of contemporary foundational modeling. In this paper, we lay the groundwork for efficient foundational modeling by proposing SAMSA - SAMpling-Self-Attention, a context-aware linear complexity self-attention mechanism that works well on multiple data modalities. Our mechanism is based on a differentiable sampling without replacement method we discovered. This enables the self-attention module to attend to the most important token set, where the importance is defined by data. Moreover, as differentiability is not needed in inference, the sparse formulation of our method costs little time overhead, further lowering computational costs. In short, SAMSA achieved competitive or even SOTA results on many benchmarks, while being faster in inference, compared to other very specialized models. Against full self-attention, real inference time significantly decreases while performance ranges from negligible degradation to outperformance. We release our source code in the repository: https://github.com/HySonLab/SAMSA
$\dagger$\$$\dagger$\$footnotetext: Correspondence to thy@uab.edu
1 Introduction
@@ 2 Related works
The scalability of transformers is hindered by the quadratic complexity of self-attention modules. Hence, numerous works have delved into sparsity (Beltagy et al., 2020; Ainslie et al., 2020; Zaheer et al., 2020; Shirzad et al., 2023), low-rank projection (Wang et al., 2020), hashing-based method (Kitaev et al., 2020), and kernel-based methods (Choromanski et al., 2021; Katharopoulos et al., 2020) to sacrifice the expressivity of transformers for a sub-quadratic construction. Our review of contemporary efficient transformers is influenced by this survey by (Tay et al., 2022). Linformer (Wang et al., 2020) is based on the low-rank assumption of self-attention matrices, realized via projection matrices. As one of the earliest works in efficient transformers, it has flaws in both scalability and fixed sequence length. Subsequent work like Linear Transformer (Katharopoulos et al., 2020) redesigns the similarity kernel previously as a softmax of dot-product score (or any other kernels) to one made of a linear probability cage. The modification frees them from the recomputation of the multiplication of key-value matrices, making attention computational cost $O(N)$, where $N$ refers to the number of tokens. Without softmax non-linearity, the rank of those attention matrices is bounded by the number of dimensions of tokens, which loses much expressivity of transformer models. Methods like (Beltagy et al., 2020) utilize three choices of fixed-sparsity-patterns: sliding window, dilated sliding window, and global+sliding window. Similarly, (Zaheer et al., 2020) uses a combination of different sparse patterns at one time: random attention, window attention, and global attention. On graphs, (Shirzad et al., 2023) also combines different sparsity patterns but with a constraint on the random-generated one: it has to be an expander graph. The hard-coded sparsity effectively linearized the complexity of transformers and is empirically fast. However, intuitively, no pattern or finite set of patterns fits all problems; therefore, there is also a line of work dedicated to a differentiable sparsity (Hy & Kondor, 2022; Nguyen et al., 2024). (Roy et al., 2021) drew a relationship from MIPS problem (Maximum Inner Product Search) and the Nearest-Neighbor Search algorithm, when both the query and key vectors are unit vectors. With their online k-means algorithm, their method attends each query vector to every key vector belonging to the same cluster, thus, bringing the complexity down to $O(n^{1.5})$, with $n$ is the number of tokens. Practically, however, their method is not fast as it needs a k-means construction and extensive specialized implementation for sparse operators. Another perspective on efficient transformers is making better implementation of vanilla transformers (Rabe & Staats, 2021; Dao et al., 2022; Dao, 2024). (Rabe & Staats, 2021) proposed a neat trick for attention mechanism: chunking. While the mechanism is simple, they have reduced the memory cost from quadratic to sublinear, effectively lifting the computation constraint of GPU memory. (Dao et al., 2022) both reduces the frequency of data transfer between HBM and SRAM within the GPU and derives a memory efficient backward pass. This allows quadratic attention scales to the sequence length of 64000 and beyond. (Dao, 2024) is an incremental work from (Dao et al., 2022) with better work partition and reduction in non-matmul FLOPS.
@@
$H(x)=\begin{cases}0&\text{if }x<0\\ 1&\text{if }x\geq 0\end{cases}$ (1)
@@
has gradient undefined at $x=0$ and zero everywhere else. Since continuous relaxation of discrete operators and algorithms is a complex study and the sampling operator is the core of our efficient transformer formulation, we focus our literature review on differentiable sampling. To continuously relax the sampling operator, the previous literature was developed based on the relaxation of choose, sort, and top-k algorithms (Jang et al., 2017; Maddison et al., 2017; Kool et al., 2019; Xie et al., 2020; Blondel et al., 2020; Petersen et al., 2021). (Jang et al., 2017; Maddison et al., 2017) is the foundational work on parameterization of discrete categorical variable. They introduces the Gumbel-Softmax distribution that approximate categorical distribution exceptionally well. There are some differences between the two works: (Jang et al., 2017) introduces straight through estimator via application of path derivative (Bengio et al., 2013) and (Maddison et al., 2017) takes account of density in their relaxed objective function. From differentiable categorical variable, differentiable sampling with replacement can be obtained trivially by computing these stochastic categorical variable $k$-times, where $k$ is the number of samples. Differentiable sampling without replacement, however, is much trickier and is an active area of research. (Kool et al., 2019) is one of the earliest work in differentiable sampling without replacement. Their process of sampling without replacement, Gumbel Top-k, samples tokens sequentially with logit normalization, applied in generative text models as stochastic beam search. (Xie et al., 2020) sees the differentiable without replacement in another perspective: optimal transport, where loss function represents how far away is the current solution from the optimal one. (Blondel et al., 2020) is an efficient estimation of sort function differentiation through optimal transport via permutahedron projections. The projection decreases the quadratic computational cost to $O(n\log{}n)$ forward pass and $O(n)$ backward pass. (Petersen et al., 2021) takes a pre-existing discrete parallelizable sorting/top-k algorithm (e.g. odd-even sorting network, bitonic sorting network, selection network, …) and relaxes all of the control flows for gradient estimation.
3 Background
3.1 Self-Attention Mechanism
Vanilla Self Attention.
@@ 3.1 Self-Attention Mechanism
The self-attention mechanism is a core component of Transformer models for long-range dependencies. Let $\mathbf{X}=\left\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\right\}$ $\in\mathbb{R}^{n\times d}$ represent a set of tokens of $n$ $d$-dimensional vectors. The matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ all in $\mathbb{R}^{n\times d_{h}}$ representing query, key, and value vectors are created via learnable linear transformation of the input token matrix $\mathbf{X}$:
$\mathbb{Q}$ $\;=\mathbb{X}W_{Q}+\mathbf{1}_{n}b_{q}^{\top},$
$\mathbb{K}$ $\;=\mathbb{X}W_{K}+\mathbf{1}_{n}b_{k}^{\top},$
$\mathbb{V}$ $\;=\mathbb{X}W_{V}+\mathbf{1}_{n}b_{v}^{\top}.$
Then, using the query, key, and value vectors, it creates another set of tokens by aggregating those value vectors based on the dependencies defined through the attention map constructed by query and value vectors:
@@
$\operatorname{Attention}(\mathbf{X})=\operatorname{Softmax}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V},$ (2)
This mechanism allows information aggregation of very far tokens and effectively captures complex long-range relationships.
Efficient Quadratic Attention.
Computing the given $\operatorname{Attention}$ function naively yields a quadratic memory cost due to the need to store attention maps for backpropagation. However, its memory complexity can be reduced into linear complexity by checkpointing (Rabe & Staats, 2021) or algebraic manipulation of derivative function (Dao et al., 2022). Via fused kernel (Dao et al., 2022; Dao, 2024), the computational cost of self-attention is further decreased through fewer memory operations between HBM and SRAM. This allows better parallelization, making quadratic attention scales to the length of 64k and beyond.
3.2 Continuous Relaxation of Top-k Sampling
Top-k Sampling.
The top-k sampling operation selects the best set of $k$ objects out of others under a criterion, deterministic or stochastic. The criterion function is a permutation invariant function that maps a set of $n$ vectors to a real number. The following equation describes a mathematical formulation of top-k sampling:
$\operatorname{Sample}(X,k)=\operatorname{argmax}_{S}^{C(X,k)}\text{criterion}(S),$ (3)
where $X$ is the set of vector to sample from, $k$ is the number of sampled vectors needing to be sampled, and $C(N,k)$ is the set of all k-combinations of set $N$. Since the $\operatorname{argmax}$ operator has no continuity, the derivative of the sampled vectors with respect to the criterion is either zero or undefined, i.e., not useful. Therefore, it is hard to integrate into neural networks. Early work (Papernot & McDaniel, 2018) had to resort to a two-stage optimization. This circumvention, however, introduces additional complexity and worsens performance via non-aligned objectives.
Straight Through Estimator.
Before going deeper into the differentiable top-k literature, it is crucial to explain a much simpler case: $k=1$. To choose one vector from many, (Jang et al., 2017) provides a differentiable approximation of the categorical distribution via Softmax with Gumbel noise added on logits:
$\operatorname{GumbelSoftmax}(\pi)[i]\triangleq\frac{\exp\left((\pi_{i}+g_{i})/\tau\right)}{\sum_{j=1}^{n}\exp\left((\pi_{j}+g_{j})/\tau\right)},$
where $g_{i}$ are independent and identically samples from Gumbel(0,1) distribution, $\pi_{i}$ are the unnormalized logits, and $\tau$ is the temperature parameter that controls the smoothness of the distribution. By adjusting $\tau$, the Gumbel-Softmax can transition between a one-hot categorical distribution and a softer distribution, enabling gradient-based optimization. The straight-through estimator, however, uses the above soft-categorical function as a differentiable proxy for the hard-categorical function obtained by $\operatorname{argmax}$:
$\operatorname{ST-GumbelSoftmax}(\pi)[i]$$\triangleq\begin{cases}1,&\text{if }i=\operatorname{argmax}(\pi),\\ 0,&\text{otherwise},\end{cases}$
$\partial{\operatorname{ST-GumbelSoftmax}}/\partial{\pi}$ $\triangleq\partial{\operatorname{GumbelSoftmax}}/\partial{\pi}.$
A naive way to extend this to differentiable top-k (Kool et al., 2019) is to use Gumbel-Softmax to sample from the unsampled vectors $k$ times sequentially, named Gumbel-Topk. This is not parallelizable and has unreliable gradients; therefore, it cannot be used inside repeating deep learning modules, e.g. transformer layers.
4 Method
4.1 Overview
@@
Figure 1: Overview of our proposed model sampling-self-attention module SAMSA. The key-value vectors are selected via top-k importance score computed using tokens’ latent. The Gumbel-Sigmoid reparameterization provides gradients to guide the optimization process towards most important key-value pairs of vectors (left). The sampled key-value vectors are then fed into Flash Attention to attend to the query vectors (right).
@@
1:  Transform tokens to query and key-value matrices (Pre-RMSNorm)
@@
3:  $\mathbf{Q}\leftarrow\widehat{\mathbb{X}}\mathbf{W_{q}}+\mathbb{1}_{n}b_{q}^{\top}$
4:  $\mathbf{P}\leftarrow\widehat{\mathbb{X}}\mathbf{W_{kv}}+\mathbb{1}_{n}b_{kv}^{\top}$
5:  Compute multi-head importance scores
6:  $\mathbf{Z}\leftarrow\operatorname{MultiHeadImportanceScore}(\mathbf{X})$
7:  $\mathbf{Z}\leftarrow\mathbf{Z}+\operatorname{Gumbel}(0,1).\operatorname{sample()}$
8:  Concatenate with 2k support vector paddings
9:  $\mathbf{P}\leftarrow\operatorname{Concatenate}(\mathbf{P},\mathbf{P_{\text{supp}}})$
10:  $\mathbf{Z}\leftarrow\operatorname{Concatenate}(\mathbf{Z},\mathbf{Z}_{\text{supp}})$
11:  Sample sets of key and value vectors
12:  $\tilde{\mathbf{P}}\leftarrow\operatorname{MultiHeadSample}(\mathbf{Z},\mathbf{P},k)$
13:  $\tilde{\mathbf{K}},\tilde{\mathbf{V}}\leftarrow\operatorname{Split}(\tilde{\mathbf{P}})$
14:  Attend and transform tokens
15:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FlashAttention}(\mathbf{Q},\tilde{\mathbf{K}},\tilde{\mathbf{V}})$
16:  $\widehat{\mathbb{X}}\leftarrow\operatorname{RMSNorm}(\mathbf{X})$
17:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FFN}(\widehat{\mathbb{X}})$
18:  Return $\mathbf{X}$
4.2 Parallelizable Differentiable Top-k Sampling
To be a viable key-value sampler for transformers, that method needs both great parallelizability and reliable gradients for efficient model training. Here, we introduce the two methods that hold the needed properties: trivial differentiable sampling with replacement and our proposed differentiable without replacement method.